OpenStudy (anonymous):

Solve the equation: |x + 5| = 9. (Points : 4)

OpenStudy (anonymous):

{4, 14}

hartnn (hartnn):

4 is correct 14 is incorrect

OpenStudy (anonymous):

x = 4 or x= -14

hartnn (hartnn):

if |a|=b a=b or a=-b

OpenStudy (anonymous):

\[x+5 = \pm9\]

OpenStudy (anonymous):

x+5=9 or x+5=-9 x=4 or x=-14

OpenStudy (anonymous):

Therefore, we get two cases. Case 1 : x + 5 = +9 Case 2 : x+5 = -9

hartnn (hartnn):

@MandyMicheala do u understand ?

OpenStudy (anonymous):

Yeah.